OpenStudy (anonymous):

expression equal to exponent 3 and root of 250

OpenStudy (anonymous):

Are you supposed to be simplifying \(\sqrt[3]{250}\) ?

OpenStudy (anonymous):

I have to find an expression equal to it

OpenStudy (anonymous):

ok. do you how to find the prime factorization of 250?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what did you get? that's all you really need to know to solve these

OpenStudy (anonymous):

\[\sqrt[3]{250}=\sqrt[3]{2 \times 5 \times 5 \times 5}=\sqrt[3]{2 \times 5^3}=5\sqrt[3]{2}\]

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

no problem